OpenStudy (anonymous):

Elena buys several shares of stock at $28 per share and sells them at $24.50 per share. What is the percent of decrease in the price of the stock? (This is confusing to me. Help!)

OpenStudy (anonymous):

24.5/28 = 0.875 1- 0.875 = 0.125 Decrease 12.5%

OpenStudy (anonymous):

28-24.5=3.5 3.5/28=.125 .125x100=12.5 12.5 decrease

OpenStudy (hoblos):

12.5 %